OpenStudy (anonymous):

4x-18y=9 Now I know you have to get y alone so I add x to the other side. Would it be -18y= 4x+9? I know how to get to the answer, you divide the right side by 18...I'm just making sure it is written down correctly.

OpenStudy (kc_kennylau):

4x-18y=9 -18y=9-4x (subtract both side by 4x) y=-(9-4x)/18 (divide both side by -18)

undeadknight26 (undeadknight26):

I thought it was: 4x-18y=9 4(0) - 18y = 9 18y = 9 /18 /18 y = .5 but now i see it makes no sense...

OpenStudy (anonymous):

So then would that make it y=-2/9 + 1/2?

OpenStudy (anonymous):

oops I meant y=-2/9x+1/2

OpenStudy (kc_kennylau):

2/9x-1/2

OpenStudy (anonymous):

Oh,, okay. Thanks!

OpenStudy (kc_kennylau):

no problem :)